The number of cars passing a motorway at any time t hours is given by the

N = 1/3t^3-5/5t^2-6t+980

find the time of which maximum number of cars passing the motorway and hence find the number of cars passing at that time.
1
Expert's answer
2020-10-09T13:39:16-0400

"N = \frac{1}{3}t^3 - \frac{5}{2}t^2 -6t +980"

Find first derivative:

"N^\prime= t^2 -5t-6"

Then

"t^2 -5t -6=0\\
t_1 + t_2 = 5\\ t_1*t_2 =-6\\
t_1 = 6\\
t_2 = -1"

Find second derivative:

"N^{\prime\prime} = 2t -5"

And find "N(-1):"

"N(-1) = 2*(-1) - 7 < 0" , so "t= -1" - maximum

But time is supposed to be nonnegative, so choose the non-negative point closest to the maximum - 0.

And number of cars at that time:

"N(0) = 980"